a is directly proportionatal to b, a=14 and b=42. Write a direct proportion equation that relate a and b.
constant of proportionality
b/a = 3/1
direct proportion equation
b/a=3/1
b = 3a

Answer by Cromlix(4381)

Hi there,
Almost right.
a = kb
Constant of proportionality
k = a/b
k = 14/42
k = 1/3
Direct proportion equation:-
a = 1/3b
Hope this helps :-)